People visiting video rental stores often rent more than one DVD at a time. The probability distribution for DVD rentals per customer at Video To Go is given in the table below. There is a five-video limit per customer at this store, so nobody ever rents more than five DVDs.
x P(x)
0 0.05
1 0.52
2 0.23
3
4 0.07
5 0.04
(a) Find the probability that a customer rents three DVDs. (Enter an exact number as an integer, fraction, or decimal.)
(b) Find the probability that a customer rents at least four DVDs. (Enter an exact number as an integer, fraction, or decimal.)
(c) Find the probability that a customer rents at most two DVDs. (Enter an exact number as an integer, fraction, or decimal.)

Answer by ikleyn(52879)

.
(a)  P(3) is the complement of the sum P(0) + P(1) + P(2) + P(4) + P(5).

     So, calculate P =  P(0) + P(1) + P(2) + P(4) + P(5) using given data.

     then find P(3) = 1 - P.



(b)  P(at least 4 DVDs) = P(4) + P(5).  Calculate using given data.


(c)  P(at most 2 DVDs) = P(0) + P(1) + P(2).  Calculate using given data.

Happy calculations !